Brandon jogs around a field that has a 490.-m perimeter in 67.0 s. Answer the following questions. Use the appropriate units: m for meter, s for second, and m/s for meter per second (a) The distance of travel d Submit Answer Tries 0/10 (b) The time of travel is t- Submit Answer Tries /10 (c) Brandons average speed is Vavg Submit Anwer Tries 0/10 (d) If Brandon arrives at the same starting position, his displacement is Ad Submit Answer Tries 0/10 (e) And his average velocity is vavg Ad/At Submit Anwer Tries 0/10 (f) Find the distance of travel if he keeps going at the same speed for another 67.0 s Submit Answer Tries /10 (g) Find how much time Brandon needs to complete a 5600-m distance if he maintains the same speed Submit Answer Tries 0/10
Answer #1

a) Distance euqal to perimeter = 490m

b) Time of travel is total time taken to cover the perimeter = 67 sec.

c) Average speed = Total distance /time = 490/67= 7.313 m/sec

d) displacement is shortest distance between two points . ending point - starting point .If both are same so net displacement is zero = 0m .

e) Average velocity = total  displacement /time = 0/67 =0 m/s

f) distance =speed * time = 7.313* 67=490 m

g)time = distance /speed= 5600/7.313=683.713 sec